OpenStudy (anonymous):

f(x)=tan^-1x then f '(1) =?

OpenStudy (anonymous):

use the formula for d/dx(tan^-1 x)=1/(1+x^2) now f '(1)=1/(1+1^2)=1/2 answ

OpenStudy (anonymous):

do you have the formula there?

OpenStudy (anonymous):

\[3^(sin2x +\cos ^{2}x) +3^(3-\sin2x-\cos^2x)=28\]show that either cos x = 0 or tan x = –1.